How do you do x+19 equals 5x+5 using bisects in geometry
Answer by farohw(175)


I believe there might be some information missing in your query???? An angle bisector divides an angle into two congruent or equal angles.

Suppose m ∠ 1 = x + 19 and m ∠ 2 = 5x + 5

To find x, simply solve the equation:

x + 19 = 5x + 5

19 - 5 = 5x - x

14 = 4x

14/4 = 4x/4

x = 7/2

Putting x = 7/2 back into the original equation,

x + 19 = 5x + 5 --> (7/2) + 19 = 5(7/2) + 5 --> 45/2 = 45/2

Therefore, each side is 22.5˚ since a 22.5˚ angle can be obtained by bisecting a 45˚ angle.